# Parametric Graph Representations in the Era of Foundation Models: A Survey and Position

### v1

@@ 2.2 High-Order Macroscopic Laws
Above discoveries are based on the node-level connections (i.e., low-order connections), then several researchers start the investigation based on the group activities, for example, motifs [40], simplices [6] and hyperedges [10, 29]. Motif is defined as a subgraph induced by a sequence of selected temporal edges in [40], where the authors discovered that different domain networks have significantly different numbers of similar motifs, and different motifs usually occur at different time. Similar laws are also discovered in [6] that the authors study 19 graph data sets from domains like biology, medicine, social networks, and the web, to characterize how high-order structure emerges and differs in different domains. They discovered that the higher-order Egonet features can discriminate the domain of the graph, and the probability of simplicial closure events typically increases with additional edges or tie strength.
In hypergraphs, each hyperedge could connect an arbitrary number of nodes, rather than two [10], where the authors found that real-world static hypergraphs obey the following properties: (1) Giant Connected Components, that there is a connected component comprising a large proportion of nodes, and this proportion is significantly larger than that of the second-largest connected component. (2) Heavy-Tailed Degree Distributions, that high-degree nodes are more likely to form new links. (3) Small Effective Diameters, that most connected pairs can be reached by a small distance (4) High Clustering Coefficients, that the global average of local clustering coefficient is high. (5) Skewed Singularvalue Distributions, that the singular-value distribution is usually heavy-tailed. Later, the evolution of real-word hypergraphs is investigated in [29], and the following laws are discovered.
@@
In heterogeneous information networks (where nodes and edges can have multiple types), the power law distribution is also discovered [47]. For example, for the triplet "author-paper-venue" (i.e., A-P-V), the number of authors is power law distributed w.r.t the number of A-P-V instances composed by an author.
@@ 3.1 Low-Order Microscopic Laws
In [34], the authors view temporal graphs in a three-fold process, i.e., node arrival (determining how many nodes will be added), edge initiation (how many edges will be added), and edge destination (where are each edge will be added). They ignore the deletion of nodes and edges, and they assign variables (models) to parameterize this process.
@@
where $a(u)$ stand for the age of node $u$, $t_{k}(u)$ is the time when node $u$ links its $k^{th}$ edge, $d_{t}(u)$ denote the degree of node $u$ at time $t$, and $d(u)=d_{T}(u)$. $T$ is the final timestamp of the data.
@@
where $\delta_{u}(d)$ records the time gap between the current time and the time when that node emits its last edge. Finding the node arrival is a regression process in [34], for example, in Flickr graph $N(t)=exp(0.25t)$, and $N(t)=3900t^{2}+76000t-130000$ in LinkedIn graph.
In [52, 41], the selection of edge attachment gets flourished where the authors propose several variants of edge attachment models for preserving the graph properties. With respect to the triangle closure phenomenon, several in-depth researches follow up. For example, in [24], researchers found that (1) the stronger the third tie (the interaction frequency of the closed edge) is, the less likely the first two ties are weakened; (2) when the stronger the first two ties are, the more likely they are weakened.
@@ 3.2 High-Order Microscopic Laws
Hypergraph ego-network [9] is a structure defined to model the high-order interactions involving an individual node. The star ego-network $T(u)$ is defined as follows.
@@
where S is the set of all hyperedges (or simplices). Also, in  [9], there are other hypergraph ego-networks, like radial ego-network $R(u)$ and contracted ego-network $C(u)$. The relationship between them is as follows.
@@
In [9], authors observe that contiguous hyperedges (simplices) in an ego-network tend to have relatively large interactions with each other, which suggests that temporally adjacent high-order interactions have high similarity, i.e., the same nodes tend to appear in neighboring simplices.
In [57], authors try to model the temporal graph growth in terms of motif evolution activities. In brief, this paper investigates how many motifs change and what are the exact motif types in each time interval and fits the arrival rate parameter of each type of motif against the whole observed temporal graph.
4 Some New Observation Space and Newly Discovered Graph Parameters
4.1 New Different Spaces
In [14], the power law is revisited based on the eigendecomposition and singular value decomposition to provide guidance on the presence of power laws in terms of the degree distribution, singular value (of adjacency matrix) distribution, and the eigenvalue (of Laplacian matrix) distribution. The authors [14] discovered that (1) degree distribution, singular value distribution, and eigenvalue distribution follow power law distribution in many real-world networks they collected; (2) and a significant power law distribution of degrees usually indicates power law distributed singular values and power law distributed eigenvalues with a high probability.
4.2 New Parameters
The local closure coefficient [54] is defined as the fraction of length-2 paths (wedges) emanating from the head node (of the wedge) that induce a triangle, i.e., starting from a seed node of a wedge, how many wedges are closed. According to [54], features extracted within the constraints of the local closure coefficient can improve the link prediction accuracy. The local closure efficient of node $u$ is defined as follows.
$H(u)=\frac{2T(u)}{W^{h}(u)}$
The density of states (or spectral density) [11] is defined as follows.
@@ 5.1 Graph Generation
If not all, in most of graph law studies [33, 56, 10, 29, 34, 41, 57], after the law (i.e., evolution pattern) is discovered, a follow-up action is to propose the corresponding graph generative model to test whether there is a realizable graph generator could generate graphs while preserving the discovered law in terms of graph properties. Also, graph generation tasks have impactful application scenarios like drug design and protein discovery [59].
For example, in [33], the Forest Fire model is proposed to preserve the macroscopic graph law while larges preserve the discovered evolution pattern.
@@
In [34], authors combine the microscopic edge destination model, edge initiation model, and node arrival rate together, to model the real-world temporal network’s growth. The parameters of these three models are fitted against the partial observation. i.e., $G_{\frac{T}{2}}$, which is the half of the entire evolving graph. Then they three produce the residual part of $G_{T}^{\prime}$. Finally, the generated $G_{T}^{\prime}$ is compared with the ground truth $G_{T}$, to see if the growth pattern is fully or near fully captured by these microscopic models. The procedures are stated as follows.
@@
  • When a node wakes up, if its lifetime has not expired yet, it creates a two-hop edge using the "random-random" triangle closing model;
@@
The generated graph $G_{T}^{\prime}$ is tested based on the comparison with the ground truth $G_{T}$, in terms of degree distribution, clustering coefficient, and diameter distribution. Taking the Flickr graph for example, the generated graph is very similar to the ground truth with aforementioned metrics [34].
@@ 5.2 Link Prediction
To learn node representation vectors for predicting links between node pairs and contributing latent applications like recommender systems, CAW-N [48] is proposed by inserting causal anonymous walks (CAWs) into the representation learning process. The CAW is a sequence of time -aware adjacent nodes, the authors claim that the extracted CAW sequence obeys the triadic closure law. To be specific, the temporal opening and closed triangles can be preserved in the extracted CAW sequence $W$. Further, to realize the inductive link prediction, CAW-N replaces the identification of each node in $W$ with the relative position information, such that the CAW sequence $W$ is transferred into anonymous $\hat{W}$. Then, the entire $\hat{W}$ is inserted into an RNN-like model and gets the embedding vector of each node, the loss function states as follows.
@@
Also, there are some related link prediction models based on the guidance of static graph laws during the representation learning process, for example, SEAL [58] and HHNE [47].
In the SEAL framework [58], for each target link, SEAL extracts a local enclosing subgraph around it, and uses a GNN to learn general graph structure features for link prediction. The corresponding graph parameters include but are not limited to
@@
5.3 Natural Language Processing
To obtain the semantic representation vector of each word in the corpus, GloVe [42] is proposed, which has been considered as one of the most popular word embedding models. GloVe utilizes the power law distribution constraint during the representation learning process. $X_{ij}$ denotes the number of times that word $j$ occurs in the context of word $i$, and it follows
$X_{ij}=\frac{k}{(r_{ij})^{\alpha}}$ (10)
where $r_{ij}$ denotes the frequency rank of the word pair $i$ and $j$ in the whole corpus, and $k$ and $\alpha$ are constant hyperparameters. Then the loss function of GloVe is stated as follows
$J=\sum^{V}_{i,j}f(X_{ij})(w_{i}^{\top}w_{j}+b_{i}+b_{j}-logX_{ij})^{2}$ (11)
where $w$ is the word vector, and $b$ is the bias vector.
@@ 6 Future Directions
In this section, we would like to list several interesting research directions of graph parametric representation in modern graph research.
@@ 6.1 Graph Laws on Temporal Graphs
Discovering accurate temporal graph laws from real-world networks heavily relies on the number of networks and the size of networks (e.g., number of nodes, number of edges, and time duration). However, some of the temporal graph law studies mentioned above usually consider the number of graphs ranging from 10 to 20, when they discover the evolution pattern. The existence of time-dependent structure and feature information increases the difficulty of collecting real-world temporal graph data. To obtain robust and accurate (temporal) graph laws, we may need a considerably large amount of (temporal) network data available. Luckily, we have seen some pioneering work like TGB [26] and TUDataset [39].
@@ 6.2 Graph Laws on Heterogeneous Networks
Though many graph laws have been proposed and verified on homogeneous graphs, real-world networks are usually heterogeneous [43] and contain a large number of interacting, multi-typed components. While the existing work [47] only studied 2 datasets to propose and verify the heterogeneous graph power law, the potential exists for a transition in graph laws from homogeneous networks to heterogeneous networks, suggesting the presence of additional parameters contributing to the comprehensive information within heterogeneous networks. For example, in an academic network, the paper citation subgraph and the author collaboration subgraph may have their own subgraph laws which affect other subgraphs’ laws. Furthermore, Knowledge graphs, as a special group of heterogeneous networks, have not yet attracted much attention from the research community to study their laws.
@@ 6.4 Taxonomy of Graph Laws
After we discovered many graph laws, is there any taxonomy or hierarchy of those? For example, graph law A stands in the superclass of graph law B, and when we preserve graph law A during the representation, we actually have already preserved graph law B. For example, there is a hierarchy of different computer vision tasks, recently discovered [55]. And corresponding research on graph law development seems like a promising direction.
@@ 6.5 Domain-Specific Graph Laws
Since graphs serve as general data representations with extreme diversity, it is challenging to find universal graph laws that fit all graph domains because each domain may be internally different from another [60]. In fact, in many cases, we have prior knowledge about the domain of a graph, which can be a social network, a protein network, or a transportation network. Thus, it is possible to study the domain-specific graph laws that work well on only a portion of graphs and then apply the graph laws only on those graphs.
@@ 6.6 Graph Laws with LLMs
In the background of large language models (LLMs) developments, an interesting question attracts much research interest nowadays, i.e., can LLMs replace GNNs as the backbone model for graphs? To answer this question, many recent works show the great efforts [25, 53, 21], where the key point is how to represent the structural information as the input for LLMs.
For example, Instruct-GLM [53] follows the manner of instruction tuning and makes the template $\mathcal{T}$ of a 2-hop connection for a central node $v$ as follows.
@@
As discussed above, the topological information (e.g., 1-hop or 2-hop connections) can serve as external modality information to contribute to (e.g., through prompting) the reasoning ability of large language models (LLMs) [25] and achieve state-of-the-art on low-order tasks like node classification and link prediction.
Therefore, a natural question can be asked, i.e., instead of inputting local topological information to LLMs, how can we bring global topological information for LLMs to understand and make inferences for high-order tasks like graph classification, graph matching, and graph alignment? To the best of our knowledge, corresponding research still remains nascent but has great potential. Finding a proper graph parametric representation in a macroscopic way may be a viable solution for LLMs to comprehend graph-level information.
@@ 7 Conclusion
Within the survey, we first review the concepts and developing progress of graph parametric representations (i.e., graph laws) from different perspectives like microscope and microscope, low-order and high-order connections, and static and temporal graphs. We then discuss various real-world application tasks that can benefit the study of graph parametric representations. Finally, we envision the latent challenges and opportunities of graph parametric representations in modern graph research with several interesting and possible future directions.

### v2

@@ 2.2 High-Order Macroscopic Laws
Above discoveries are based on the node-level connections (i.e., low-order connections), then several researchers start the investigation based on the group activities, for example, motifs Paranjape et al. (2017), simplices Benson et al. (2018) and hyperedges Do et al. (2020); Kook et al. (2020). Motif is defined as a subgraph induced by a sequence of selected temporal edges in Paranjape et al. (2017), where the authors discovered that different domain networks have significantly different numbers of similar motifs, and different motifs usually occur at different time. Similar laws are also discovered in Benson et al. (2018) that the authors study 19 graph data sets from domains like biology, medicine, social networks, and the web, to characterize how high-order structure emerges and differs in different domains. They discovered that the higher-order Egonet features can discriminate the domain of the graph, and the probability of simplicial closure events typically increases with additional edges or tie strength.
In hypergraphs, each hyperedge could connect an arbitrary number of nodes, rather than two Do et al. (2020), where the authors found that real-world static hypergraphs obey the following properties: (1) Giant Connected Components, that there is a connected component comprising a large proportion of nodes, and this proportion is significantly larger than that of the second-largest connected component. (2) Heavy-Tailed Degree Distributions, that high-degree nodes are more likely to form new links. (3) Small Effective Diameters, that most connected pairs can be reached by a small distance (4) High Clustering Coefficients, that the global average of local clustering coefficient is high. (5) Skewed Singularvalue Distributions, that the singular-value distribution is usually heavy-tailed. Later, the evolution of real-word hypergraphs is investigated in Kook et al. (2020), and the following laws are discovered.
@@
In heterogeneous information networks (where nodes and edges can have multiple types), the power law distribution is also discovered Wang et al. (2019). For example, for the triplet ”author-paper-venue” (i.e., A-P-V), the number of authors is power law distributed w.r.t the number of A-P-V instances composed by an author.
@@ 3.1 Low-Order Microscopic Laws
In Leskovec et al. (2008), the authors view temporal graphs in a three-fold process, i.e., node arrival (determining how many nodes will be added), edge initiation (how many edges will be added), and edge destination (where are each edge will be added). They ignore the deletion of nodes and edges, and they assign variables (models) to parameterize this process.
@@
where $a(u)$ stands for the age of node $u$, $t_{k}(u)$ is the time when node $u$ links its $k^{th}$ edge, $d_{t}(u)$ denote the degree of node $u$ at time $t$, and $d(u)=d_{T}(u)$. $T$ is the final timestamp of the data.
@@
where $\delta_{u}(d)$ records the time gap between the current time and the time when that node emits its last edge. Finding the node arrival is a regression process in Leskovec et al. (2008), for example, in Flickr graph $N(t)=exp(0.25t)$, and $N(t)=3900t^{2}+76000t-130000$ in LinkedIn graph.
In Yang et al. (2013); Park and Kim (2018), the selection of edge attachment gets flourished where the authors propose several variants of edge attachment models for preserving the graph properties. With respect to the triangle closure phenomenon, several in-depth researches follow up. For example, in Huang et al. (2018), researchers found that (1) the stronger the third tie (the interaction frequency of the closed edge) is, the less likely the first two ties are weakened; (2) when the stronger the first two ties are, the more likely they are weakened.
@@ 3.2 High-Order Microscopic Laws
Hypergraph ego-network Comrie and Kleinberg (2021) is a structure defined to model the high-order interactions involving an individual node. The star ego-network $T(u)$ is defined as follows.
@@
where S is the set of all hyperedges (or simplices). Also, in  Comrie and Kleinberg (2021), there are other hypergraph ego-networks, like radial ego-network $R(u)$ and contracted ego-network $C(u)$. The relationship between them is as follows.
@@
In Comrie and Kleinberg (2021), authors observe that contiguous hyperedges (simplices) in an ego-network tend to have relatively large interactions with each other, which suggests that temporally adjacent high-order interactions have high similarity, i.e., the same nodes tend to appear in neighboring simplices.
In Zeno et al. (2020), authors try to model the temporal graph growth in terms of motif evolution activities. In brief, this paper investigates how many motifs change and what are the exact motif types in each time interval and fits the arrival rate parameter of each type of motif against the whole observed temporal graph.
@@ 4.1 Graph Generation
If not all, in most of graph law studies Leskovec et al. (2005); Zang et al. (2018); Do et al. (2020); Kook et al. (2020); Leskovec et al. (2008); Park and Kim (2018); Zeno et al. (2020), after the law (i.e., evolution pattern) is discovered, a follow-up action is to propose the corresponding graph generative model to test whether there is a realizable graph generator could generate graphs while preserving the discovered law in terms of graph properties. Also, graph generation tasks have impactful application scenarios like drug design and protein discovery Zhang and et al. (2023).
For example, in Leskovec et al. (2005), the Forest Fire model is proposed to preserve the macroscopic graph law while larges preserve the discovered evolution pattern.
@@
In Leskovec et al. (2008), the authors combine the microscopic edge destination model, edge initiation model, and node arrival rate together, to model the real-world temporal network’s growth. The parameters of these three models are fitted against the partial observation. i.e., $G_{\frac{T}{2}}$, which is the half of the entire evolving graph. Then they three produce the residual part of $G_{T}^{\prime}$. Finally, the generated $G_{T}^{\prime}$ is compared with the ground truth $G_{T}$, to see if the growth pattern is fully or near fully captured by these microscopic models. The procedures are stated as follows.
@@
  • When a node wakes up, if its lifetime has not expired yet, it creates a two-hop edge using the ”random-random” triangle closing model;
@@
The generated graph $G_{T}^{\prime}$ is tested based on the comparison with the ground truth $G_{T}$, in terms of degree distribution, clustering coefficient, and diameter distribution. Taking the Flickr graph for example, the generated graph is very similar to the ground truth with aforementioned metrics Leskovec et al. (2008).
@@ 4.2 Link Prediction
To learn node representation vectors for predicting links between node pairs and contributing latent applications like recommender systems, CAW-N Wang et al. (2021) is proposed by inserting causal anonymous walks (CAWs) into the representation learning process. The CAW is a sequence of time -aware adjacent nodes, the authors claim that the extracted CAW sequence obeys the triadic closure law. To be specific, the temporal opening and closed triangles can be preserved in the extracted CAW sequence $W$. Further, to realize the inductive link prediction, CAW-N replaces the identification of each node in $W$ with the relative position information, such that the CAW sequence $W$ is transferred into anonymous $\hat{W}$. Then, the entire $\hat{W}$ is inserted into an RNN-like model and gets the embedding vector of each node, the loss function states as follows.
@@
Also, there are some related link prediction models based on the guidance of static graph laws during the representation learning process, for example, SEAL Zhang and Chen (2018) and HHNE Wang et al. (2019).
In the SEAL framework Zhang and Chen (2018), for each target link, SEAL extracts a local enclosing subgraph around it, and uses a GNN to learn general graph structure features for link prediction. The corresponding graph parameters include but are not limited to
@@
5 Some New Observation Space and Newly Discovered Graph Parameters
5.1 New Different Spaces
In Eikmeier and Gleich (2017), the power law is revisited based on the eigendecomposition and singular value decomposition to provide guidance on the presence of power laws in terms of the degree distribution, singular value (of adjacency matrix) distribution, and the eigenvalue (of Laplacian matrix) distribution. The authors Eikmeier and Gleich (2017) discovered that (1) degree distribution, singular value distribution, and eigenvalue distribution follow power law distribution in many real-world networks they collected; (2) and a significant power law distribution of degrees usually indicates power law distributed singular values and power law distributed eigenvalues with a high probability.
5.2 New Parameters
The local closure coefficient Yin et al. (2019) is defined as the fraction of length-2 paths (wedges) emanating from the head node (of the wedge) that induce a triangle, i.e., starting from a seed node of a wedge, how many wedges are closed. According to Yin et al. (2019), features extracted within the constraints of the local closure coefficient can improve the link prediction accuracy. The local closure efficient of node $u$ is defined as follows.
$H(u)=\frac{2T(u)}{W^{h}(u)}$
The density of states (or spectral density) Dong et al. (2019) is defined as follows.
6 Related Work
To the best of our knowledge, there are only a few survey papers on graph laws, with none published after 2022, marking the beginning of the foundation model era. A 2006 survey Chakrabarti and Faloutsos (2006) primarily focused on graph laws for mining patterns, discussing the Densification Law and Shrinking Law. In 2016, another survey Drakopoulosa et al. (2016) shifted its focus towards the generation of large graphs using different graph modeling methods, including the Erdos-Renyi model, Watts-Strogatz model, and Albert-Barabasi model. More recently, in 2019, the authors in Drobyshevskiy and Turdakov (2020) offered a broader perspective on random graph modeling, covering generative, feature-driven, and domain-specific approaches. In contrast to these earlier surveys, which were published before the advent of graph neural networks and prior to the discovery of several significant graph laws Yin et al. (2019); Dong et al. (2019); Zeno et al. (2020); Do et al. (2020); Kook et al. (2020); Comrie and Kleinberg (2021); Wang et al. (2019), our work represents the first survey to explore the potential of graph laws in the context of foundation models. We emphasize how graph laws can address domain inconsistencies across different graph data types and contribute to multimodal representation learning. Additionally, this survey is the first to offer an overview of high-order graph laws and heterogeneous graph laws, marking a novel contribution to the literature.
@@ 7 Future Directions
Here, we list several interesting research directions of graph parametric representation in modern graph research.
@@ 7.1 Graph Laws on Temporal Graphs
Discovering accurate temporal graph laws from real-world networks heavily relies on the number of networks and the size of networks (e.g., number of nodes, number of edges, and time duration). However, some of the temporal graph law studies mentioned above usually consider the number of graphs ranging from 10 to 20, when they discover the evolution pattern. The existence of time-dependent structure and feature information increases the difficulty of collecting real-world temporal graph data. To obtain robust and accurate (temporal) graph laws, we may need a considerably large amount of (temporal) network data available. Luckily, we have seen some pioneering work like TGB Huang et al. (2023b) and TUDataset Morris et al. (2020).
@@ 7.2 Graph Laws on Heterogeneous Networks
Though many graph laws have been proposed and verified on homogeneous graphs, real-world networks are usually heterogeneous Shi et al. (2017) and contain a large number of interacting, multi-typed components. While the existing work Wang et al. (2019) only studied 2 datasets to propose and verify the heterogeneous graph power law, the potential exists for a transition in graph laws from homogeneous networks to heterogeneous networks, suggesting the presence of additional parameters contributing to the comprehensive information within heterogeneous networks. For example, in an academic network, the paper citation subgraph and the author collaboration subgraph may have their own subgraph laws which affect other subgraphs’ laws. Furthermore, Knowledge graphs, as a special group of heterogeneous networks, have not yet attracted much attention from the research community to study their laws.
@@ 7.4 Taxonomy of Graph Laws
After we discovered many graph laws, is there any taxonomy or hierarchy of those? For example, graph law A stands in the superclass of graph law B, and when we preserve graph law A during the representation, we actually have already preserved graph law B. For example, there is a hierarchy of different computer vision tasks, recently discovered Zamir et al. (2018). And corresponding research on graph law development seems like a promising direction.
@@ 7.5 Domain-Specific Graph Laws
Since graphs serve as general data representations with extreme diversity, it is challenging to find universal graph laws that fit all graph domains because each domain may be internally different from another Zhang et al. (2023). In fact, in many cases, we have prior knowledge about the domain of a graph, which can be a social network, a protein network, or a transportation network. Thus, it is possible to study the domain-specific graph laws that work well on only a portion of graphs and then apply the graph laws only on those graphs.
@@ 7.6 Graph Laws with LLMs
In the background of large language models (LLMs) developments, an interesting question attracts much research interest nowadays, i.e., can LLMs replace GNNs as the backbone model for graphs? To answer this question, many recent works show the great efforts Huang et al. (2023a); Ye et al. (2024); He et al. (2024a), where the key point is how to represent the structural information as the input for LLMs.
For example, Instruct-GLM Ye et al. (2024) follows the manner of instruction tuning and makes the template $\mathcal{T}$ of a 2-hop connection for a central node $v$ as follows.
@@
As discussed above, the topological information (e.g., 1-hop or 2-hop connections) can serve as external modality information to contribute to (e.g., through prompting) the reasoning ability of large language models (LLMs) Huang et al. (2023a) and achieve state-of-the-art on low-order tasks like node classification and link prediction.
@@ 8 Conclusion
Motivated by LLM’s need to understand graphs, we first review the concepts and development progress of graph parametric representations (i.e., graph laws) from different perspectives like macroscope and microscope, low-order and high-order connections, and static and temporal graphs. We then discuss various real-world application tasks that can benefit the study of graph parametric representations. Finally, we envision the latent challenges and opportunities of graph parametric representations in modern graph research with several interesting and possible future directions.